# mywords.dic

### v1

@@ Abstract
Reinforcement Learning from Verifiable Rewards (RLVR) has emerged as a promising framework for enhancing the reasoning capabilities of large language models. However, existing approaches such as GRPO often suffer from zero gradients. This problem arises primarily due to fixed clipping bounds for token-level probability ratios and the standardization of identical rewards, which can lead to ineffective gradient updates and underutilization of generated responses. In this work, we propose Dynamic Clipping Policy Optimization (DCPO), which introduces a dynamic clipping strategy that adaptively adjusts the clipping bounds based on token-specific prior probabilities to enhance token-level exploration, and a smooth advantage standardization technique that standardizes rewards across cumulative training steps to improve the response-level effective utilization of generated responses. DCPO achieved state-of-the-art performance on four benchmarks based on four different models. In particular, DCPO achieved an Avg@1 of 46.7 under greedy decoding and an Avg@32 of 38.8 under 32 times sampling on the AIME24 benchmark, surpassing both DAPO (36.7/31.6) and GRPO (36.7/32.1) on the Qwen2.5-Math-7B model. On the AIME25 benchmark based on Qwen2.5-14B, DCPO achieves a performance of (23.3/19.0), surpassing GRPO (13.3/10.5) and DAPO (20.0/15.3). Furthermore, DCPO achieved an average 28% improvement in the nonzero advantage over GRPO in four models, doubled the training efficiency over DAPO, and significantly reduced the token clipping ratio by an order of magnitude compared to both GRPO and DAPO, while achieving superior performance. These results highlight DCPO’s effectiveness in leveraging generated data more efficiently for reinforcement learning in large language models.
@@ 1 Introduction
To support RLVR, the GRPO (DeepSeek-AI et al., 2024; Liu et al., 2025; Yuan et al., 2025) family of algorithms was developed. Although the robust performance of GRPO has been extensively validated in numerous studies, recent research, such as DAPO (Yu et al., 2025), has identified critical limitations. Specifically, the GRPO suffers from entropy collapse, characterized by the advantage values decaying to zero during later training stages. Furthermore, inappropriate weight-clipping can discard rarely sampled learning signals, thus overly limiting the diversity of model exploration.
@@
In this work, we introduce Dynamic Clipping Policy Optimization (DCPO), a novel approach designed to address the key limitations of previous methods. First, we introduce a smoothing technique that standardizes advantages across cumulative steps to address the gradient descent problem caused by zeroing out the advantages of responses with the same reward. This method aggregates the reward distribution of all cumulatively generated responses with the current generated distribution, allowing more efficient response-level exploration(see Section 2). Second, we propose a dynamic clipping method that adaptively adjusts the clipping bounds based on the old token-specific probabilities to mitigate the drawbacks of the fixed clipping bounds, rather than setting different fixed upper and lower clipping bounds. Specifically, tokens with lower prior probabilities are granted wider clipping bounds, allowing more space for token-level exploration(see Section 2 and Appendix A.6). In addition, DCPO only computes the mean loss across tokens within each individual response instead of averaging the loss over an entire batch, which preserves the relative advantage structure among responses to the same prompt and avoids the dilution effect introduced by batch-level averaging.
Finally, we merged Math-DAPO-17k (Yu et al., 2025) with levels 3-5 of the MATH dataset (Liu et al., 2025) as the training dataset and selected four common mathematical reasoning benchmarks to evaluate performance. We conducted experiments on four different models and used GRPO and DAPO as baselines. The experimental results show that DCPO consistently achieved the superior preference, with particularly large gains on the challenging AIME benchmark: on AIME24-Avg@32 with 32-time sampling, DCPO-7B scored 38.8 vs. 32.1 (GRPO) and 31.6 (DAPO), and on AIME25-Avg@32, DCPO-14B reached 19.0 vs. 10.5 and 15.3. Moreover, DCPO increased the average nonzero advantage ratio by 28% over GRPO, doubled training efficiency over DAPO, and reduced token clipping by an order of magnitude compared to all baselines, while delivering superior performance.
For preliminary details, including loss and advantage calculation methods for previous algorithms such as GRPO and DAPO, see the Appendix A.2.
@@ 2 Dynamic-Adaptive Clipping Bounds for RL(DAC)
However, a symmetric fixed clipping bound constrains the model’s exploration of the diversity space, particularly in regions where the old policy assigns low probability mass. In these scenarios, the absolute bound for decreasing probabilities is inherently limited, restricting the capacity for meaningful policy updates. Although methods like DAPO attempt to address this issue with asymmetric clipping bounds, the fundamental limitation persists.
@@ 3 Smooth Advantage Standardization(SAS)
In previous works, such as GRPO and DAPO, the advantage $\hat{A}^{i}_{new,j}$} of the $j$-th response is the standardized result of only taking the rewards $R_{j=1,\dots,G}^{i}$ for the same prompt at the $i$-th step, as Equation 11 in the Appendix A.3. However, this approach can lead to several issues: First, when randomness in response sampling causes all rewards to be the same at a given step, the advantage becomes zero, preventing the prompt from contributing to parameter updates despite potentially valuable differences in reasoning trajectories. Second, randomness in high-entropy sampling can yield highly skewed label counts, causing large fluctuations in standardized advantage values across steps, even reversing signs, thus destabilizing training.
@@
$\hat{SA}^{i}_{new,j}$} places more weight on the current distribution as training progresses, reflecting a preference for adapting to the latest policy outputs. Conversely, $\hat{SA}^{i}_{total,j}$} places more weight on the overall cumulative distribution, viewing the deviation of the current step as a fine-tuning adjustment within the broader distribution. Finally, our final advantage $\hat{A}^{i}_{j}$ is defined as the smoothed advantage with the smaller absolute value to reduce the impact of the respective fluctuations of cumulative standardization and standardization of the current step on training stability, defined as Equation 7
@@ 4 Experimental Setup
We used GRPO and DAPO as baselines and added GSPO to the ablation baselines. We set a symmetric clipping threshold of $\epsilon=0.2$ for GRPO and lower/upper clipping thresholds $(0.2,0.28)$ with an additional 512-token soft-punish cache for DAPO, lower/upper clipping thresholds (3e4,4e4)(3e-4,4e-4) for GSPO used in ablation studies, and low/high clipping thresholds $(0.16,0.2)$ for DCPO, preserving the same clipping points $(q(x),r(x))=(\frac{1}{1+\epsilon},1+\epsilon)$ and $(q(x),r(x))=(1,1-\epsilon)$ as GRPO (Figure 4(d)), and set the maximum probability ratio to 10 for both positive and negative advantages, different from DAPO. The other common settings, such as the codebase and training hyperparameters, are described in the Appendix A.9.
@@ 5.1 Main Results
Table 1 summarizes the performance of various preference optimization methods, including GRPO, DAPO, and our proposed DCPO, on four math reasoning benchmarks: MATH500, AMC23, AIME24, and AIME25. We report both greedy decoding results (Avg@1) and sampling-based decoding performance (Avg@32), providing a comprehensive evaluation of the robustness and generalization of each method. Across all model scales, DCPO consistently demonstrates superior or comparable performance relative to GRPO and DAPO, indicating its effectiveness as a general-purpose preference optimization technique in reinforcement learning of LLMs.
@@
On the Qwen2.5-Math-1.5B-Instruct, DCPO achieves the highest average scores (47.2/32.8), exceeding GRPO (+1.2/+0.8) and DAPO (+0.7/+0.4). On Qwen2.5-3B, DCPO yields (37.6/22.7), exceeding GRPO (+1.3/+1.7) and performing competitively with DAPO, with a notable improvement in Avg@1 (+1.0) and only a marginal gap in Avg@32 (-0.4). For Qwen2.5-Math-7B, DCPO attains (57.1/45.2), surpassing GRPO (+4.0/+3.6) and DAPO (+3.3/+2.8). On the largest tested model, Qwen2.5-14B, DCPO delivers a strong result of (53.2/39.0), offering substantial improvements over GRPO (+6.5/+7.7) and outperforming DAPO by (+1.3/+0.1) in Avg@1 and Avg@32. These results highlight DCPO’s robust improvements across different model capacities and its superior alignment to the reward signal under both greedy and stochastic decoding paradigms.
@@
A particularly notable trend emerges on the AIME24 and AIME25 datasets, high-difficulty competition math benchmarks designed to stress-test a model’s reasoning ability. Under Avg@32, which evaluates the model’s ability to generate correct responses under high-entropy sampling, DCPO exhibits significantly enhanced robustness on the 7B and 14B scales: On AIME24-Avg@32, DCPO-7B achieves 38.8, outperforming GRPO (32.1) and DAPO (31.6). On AIME25-Avg@32, DCPO-14B reaches 19.0, exceeding GRPO (10.5) and DAPO (15.3). These gains suggest that DCPO not only improves average case performance, but also enhances the diversity and reliability of correct generations, which is particularly critical under sampling-based decoding where beam search or temperature-driven variability is present. The method’s ability to optimize both reward fidelity and exploration under high entropy decoding highlights its practical value in real-world settings where multiple plausible outputs are sampled.
@@ 5.2 The Token Clipping Ratio (TCR)
In GRPO training, the TCR exhibits divergent trends based on the size of model. For smaller models (Qwen2.5-Math-1.5B and Qwen2.5-3B), the TCR increases with training steps, while for larger models (Qwen2.5-Math-7B and Qwen2.5-14B), it gradually decreases. In contrast, DAPO consistently shows an upward trajectory in the TCR for all models. This indicates that as DAPO training progresses, an increasing number of tokens are clipped, leading to a higher proportion of policy updates based on partial or truncated responses. However, both GRPO and DAPO exhibit significant fluctuations and occasional abnormal spikes in the TCR, which may introduce instability into the optimization process and potentially lead to entropy collapses.
@@
Previous work (Wang et al., 2025) demonstrated that high-entropy tokens, which are precisely the low-probability tokens, are the primary drivers of the model’s emergence of reasoning capabilities. Because entropy and probability are inversely related, the rarer a token, the more informational content it carries. Therefore, providing a broader admissible update range for such low-probability tokens enlarges the effective exploration space of the model, facilitating the acquisition of diverse reasoning capabilities.
@@ 5.3 The Response Utilization Ratio (RUR)
During RLVR training, a policy update is contingent on a non-zero gradient for a given response. We quantify the efficiency of this process with a metric defined as the Response Utilization Ratio (RUR), which is the percentage of non-zero advantage of generated responses that participate in the policy update.
@@
Section 5.3 summarizes the overall average RUR for responses generated during GRPO training is low, with an overall mean of only 44.6% across four different models. This indicates that more than half of generated responses are discarded, highlighting the significant inefficiency in response utilization during model diversity exploration. In contrast, our proposed DCPO method achieves an average RUR of approximately 70% after the first epoch. This represents a substantial improvement, with an absolute increase of 28% and a relative improvement of 64% over the GRPO’s.
@@
model GRPO DCPO
Qwen2.5-Math-1.5B-Instruct 45.6% 67.1%
Qwen2.5-3B 48.3% 74.3%
Qwen2.5-Math-7B 37.4% 73.2%
Qwen2.5-14B 43.9% 72.4%
Average 43.8% 71.8%
@@ 5.4 Ablation Studies
We conducted an ablation study on Qwen2.5-Math-7B to assess the contribution of each component in DCPO, using Avg@32 as the evaluation metric. This metric highlights the robustness and stability of the learned policy distribution. To ensure fairness, each experiment modifies a single component of the baseline GRPO framework while keeping all other settings identical. In addition, we added GSPO (Zheng et al., 2025) as an additional baseline. GSPO uses sequence-level clipping instead of token-level clipping to reduce the proportion of high-entropy responses participating in model updates, resulting in most responses with non-zero advantage being wasted. The learning curves among 20 to 400 steps are shown in Figure 3.
@@ A.3 Advantage Calculation
In previous works, including GRPO and DAPO, the advantage $\hat{A}_{j,t}^{i}$ for the token $t$ in the response $j$ is calculated by standardizing the reward $R^{i}_{j}$ against the mean $\mu^{i}$ and standard deviation $\sigma^{i}$ of the rewards of the $G$ generated responses at the $i$-th step, as shown in  Equation 11.
@@ A.3.1 GRPO
Shao et al. (2024) proposed the Group Relative Policy Optimization(GRPO), which first samples $G$ responses for each query, assigns rewards $R$ via a rule-based reward function, and estimates token-level advantages as in Equation 11. Finally, the model parameters are updated using Equation 12, referred to as Sequence Level Mean loss (SLM) as:

### v2

@@ Abstract
Reinforcement Learning from Verifiable Rewards (RLVR) has emerged as a promising framework for enhancing the reasoning capabilities of large language models. However, existing approaches such as GRPO often suffer from zero gradients. This problem arises primarily due to fixed clipping bounds for token-level probability ratios and the standardization of identical rewards, which can lead to ineffective gradient updates and underutilization of generated responses. In this work, we propose Dynamic Clipping Policy Optimization (DCPO), which introduces a dynamic clipping strategy that adaptively adjusts clipping bounds based on token-specific prior probabilities to enhance token-level exploration, and a smooth advantage standardization technique that standardizes rewards across cumulative training steps to improve the response-level effective utilization of generated responses. DCPO achieved state-of-the-art performance on four benchmarks based on four different models. In particular, DCPO achieved an Avg@1 of 46.7 under greedy decoding and an Avg@32 of 38.8 under 32 times sampling on the AIME24 benchmark, surpassing DAPO (36.7/31.6), GRPO (36.7/32.1) and GSPO (40.0/34.9) on the Qwen2.5-Math-7B model. On the AIME25 benchmark based on Qwen2.5-14B, DCPO achieves a performance of (23.3/19.0), surpassing GRPO (13.3/10.5), DAPO (20.0/15.3) and GSPO (16.7/9.9). Furthermore, DCPO achieved an average 28% improvement in the nonzero advantage over GRPO in four models, doubled the training efficiency over DAPO, and significantly reduced the token clipping ratio by an order of magnitude compared to both GRPO and DAPO, while achieving superior performance. These results highlight DCPO’s effectiveness in leveraging generated data more efficiently for reinforcement learning in large language models.
@@ 1 Introduction
To support RLVR, the GRPO (DeepSeek-AI et al., 2024; Liu et al., 2025; Yuan et al., 2025; Zheng et al., 2025) family of algorithms was developed. Although the robust performance of GRPO has been extensively validated in numerous studies, recent research, such as DAPO (Yu et al., 2025), has identified critical limitations. Specifically, the GRPO suffers from entropy collapse, characterized by the advantage values decaying to zero during later training stages. Furthermore, inappropriate weight-clipping can discard rarely sampled learning signals, thus overly limiting the diversity of model exploration.
@@
In this work, we introduce Dynamic Clipping Policy Optimization (DCPO), a novel approach designed to address the key limitations of previous methods. First, we introduce a smoothing technique that standardizes advantages across cumulative steps to address the gradient descent problem caused by zeroing out the advantages of responses with the same reward. This method aggregates the reward distribution of all cumulatively generated responses with the current generated distribution, allowing more efficient response-level exploration(see Section 2). Second, we propose a dynamic clipping method that adaptively adjusts the clipping bounds based on the old token-specific probabilities to mitigate the drawbacks of the fixed clipping bounds, rather than setting different fixed upper and lower clipping bounds. Specifically, tokens with lower prior probabilities are granted wider clipping bounds, allowing more space for token-level exploration(see Section 2 and Appendix A.6). In addition, DCPO only computes the mean loss across tokens within each individual response instead of averaging over an entire batch, which preserves the relative advantage structure among responses to the same prompt and avoids the dilution effect of batch-level averaging.
Finally, we merged Math-DAPO-17k (Yu et al., 2025) with levels 3-5 of the MATH dataset (Liu et al., 2025) as the training dataset and selected four common mathematical reasoning benchmarks to evaluate performance. We conducted experiments on four different models and used GRPO, DAPO and (where available) GSPO as baselines. The experimental results show that DCPO consistently achieved the superior preference, with particularly large gains on the challenging AIME benchmarks: on AIME24-Avg@32 with 32-time sampling, DCPO-7B scored 38.8 vs. 32.1 (GRPO) vs. 31.6 (DAPO) vs. 34.9 (GSPO), and on AIME25-Avg@32, DCPO-14B reached 19.0 vs. 10.5 (GRPO) vs. 15.3 (DAPO) vs. 9.9 (GSPO). Moreover, DCPO increased the average nonzero advantage ratio by 28% over GRPO, doubled training efficiency over DAPO, and reduced token clipping by an order of magnitude compared to all baselines, while delivering superior performance.
For preliminary details, including loss and advantage calculation methods for previous algorithms such as GRPO, DAPO and GSPO, see the Appendix A.2.
@@ 2 Dynamic-Adaptive Clipping Bounds for RL(DAC)
However, a symmetric fixed clipping bound constrains the model’s exploration of the diversity space, particularly in regions where the old policy assigns low probability mass. In these scenarios, the absolute bound for decreasing probabilities is inherently limited, restricting the capacity for meaningful policy updates. Although methods like DAPO attempt to address this issue with asymmetric clipping bounds, the fundamental limitation persists. Furthermore, GSPO discards entire responses when the sequence-level importance ratio exceeds the asymmetric clipping bounds, wasting more informative tokens, and retaining more high-variance tokens leading to training instability.
@@ 3 Smooth Advantage Standardization(SAS)
In previous works, such as GRPO, DAPO and GSPO, the advantage $\hat{A}^{i}_{new,j}$} of the $j$-th response is the standardized result of only taking the rewards $R_{j=1,\dots,G}^{i}$ for the same prompt at the $i$-th step, as Equation 11 in the Appendix A.3. However, this approach can lead to several issues: First, when randomness in response sampling causes all rewards to be the same at a given step, the advantage becomes zero, preventing the prompt from contributing to parameter updates despite potentially valuable differences in reasoning trajectories. Second, randomness in high-entropy sampling can yield highly skewed label counts, causing large fluctuations in standardized advantage values across steps, even reversing signs, thus destabilizing training.
@@
$\hat{SA}^{i}_{new,j}$} places more weight on the current distribution as training progresses, reflecting a preference for adapting to the latest policy outputs. Conversely, $\hat{SA}^{i}_{total,j}$} places more weight on the overall cumulative distribution, viewing the deviation of the current step as a fine-tuning adjustment within the broader distribution. Finally, our final advantage $\hat{A}^{i}_{j}$ is defined as the smoothed advantage with the smaller absolute value to reduce the impact of the respective fluctuations of cumulative standardization and standardization of the current step on training stability, defined as Equation 7.
@@ 4 Experimental Setup
We used GRPO and DAPO as baselines and added GSPO to Qwen2.5-Math-7B and Qwen2.5-14B as additional baselines. The other settings, such as the different clipping thresholds $\epsilon_{*}$, codebase, and other common training hyperparameters, are described in the Appendix A.9.
@@ 5.1 Main Results
$\Delta_{GSPO}$ +6.0 +7.5/4.9 -3.3/+2.2 +6.6/+10.1 +4.2/+5.5
Table 1 summarizes the performance of various preference optimization methods, including GRPO, DAPO, GSPO, and our proposed DCPO, on four math reasoning benchmarks: MATH500, AMC23, AIME24, and AIME25. We report both greedy decoding results (Avg@1) and sampling-based decoding performance (Avg@32), providing a comprehensive evaluation of the robustness and generalization of each method. Across all model scales, DCPO consistently demonstrates superior or comparable performance relative to GRPO, DAPO and (where available) GSPO, indicating its effectiveness as a general-purpose preference optimization technique in reinforcement learning of LLMs.
@@
On the Qwen2.5-Math-1.5B-Instruct, DCPO achieves the highest average scores (47.2/32.8), exceeding GRPO (+1.2/+0.8) and DAPO (+0.7/+0.4). On Qwen2.5-3B, DCPO yields (37.6/22.7), exceeding GRPO (+1.3/+1.7) and performing competitively with DAPO, with a notable improvement in Avg@1 (+1.0) and only a marginal gap in Avg@32 (-0.4). For Qwen2.5-Math-7B, DCPO attains (57.1/45.2), surpassing GRPO (+4.0/+3.6) and DAPO (+3.3/+2.8) and GSPO (+1.9/+1.9). On the largest tested model, Qwen2.5-14B, DCPO delivers a strong result of (53.2/39.0), offering substantial improvements over GRPO (+6.5/+7.7) and outperforming DAPO by (+1.3/+0.1) and GSPO by (+4.2/+5.5) in Avg@1 and Avg@32. These results highlight DCPO’s robust improvements across different model capacities and its superior alignment to the reward signal under both greedy and stochastic decoding paradigms.
@@
A particularly notable trend emerges on the AIME24 and AIME25 datasets, high-difficulty competition math benchmarks designed to stress-test a model’s reasoning ability. Under Avg@32, which evaluates the model’s ability to generate correct responses under high-entropy sampling, DCPO exhibits significantly enhanced robustness on the 7B and 14B scales: On AIME24-Avg@32, DCPO-7B achieves 38.8, outperforming GRPO (32.1), DAPO (31.6) and GSPO (34.9). On AIME25-Avg@32, DCPO-14B reaches 19.0, exceeding GRPO (10.5), DAPO (15.3) and GSPO (9.9). These gains suggest that DCPO not only improves average case performance, but also enhances the diversity and reliability of correct generations, which is particularly critical under sampling-based decoding where beam search or temperature-driven variability is present. The method’s ability to optimize both reward fidelity and exploration under high entropy decoding highlights its practical value in real-world settings where multiple plausible outputs are sampled.
@@ 5.2 The Token Clipping Ratio (TCR)
In GRPO training, the TCR exhibits divergent trends based on the size of model. For smaller models (Qwen2.5-Math-1.5B and Qwen2.5-3B), the TCR increases with training steps, while for larger models (Qwen2.5-Math-7B and Qwen2.5-14B), it gradually decreases. In contrast, DAPO consistently shows an upward trajectory in the TCR for all models. This indicates that as DAPO training progresses, an increasing number of tokens are clipped, leading to a higher proportion of policy updates based on partial or truncated responses. However, both GRPO and DAPO exhibit significant fluctuations and occasional abnormal spikes in the TCR, which may introduce instability into the optimization process and potentially lead to entropy collapses. At the same time, the average TCR of GSPO based on Qwen2.5-Math-7B is greater than 11%, and based on Qwen2.5-14B is over 15%, which is much higher than the average TCR of token-level clipping methods and wastes more informative tokens.
@@
Previous work (Wang et al., 2025) demonstrated that high-entropy tokens, which are precisely low-probability tokens, are the primary drivers of the model’s emergence of reasoning capabilities. Because entropy and probability are inversely related, the rarer a token, the more informational content it carries. Therefore, providing a broader admissible update range for such low-probability tokens enlarges the effective exploration space of the model, facilitating the acquisition of diverse reasoning capabilities.
@@ 5.3 The Response Utilization Ratio (RUR)
During RLVR training, a policy update is contingent on a nonzero gradient for a given response. We quantify the efficiency of this process with a metric defined as the Response Utilization Ratio (RUR), which is the percentage of nonzero advantage of generated responses that participate in the policy update.
@@
Section 5.3 summarizes the overall average RUR for responses generated during GRPO training is low, with an overall mean of only 44.6% across four different models. This indicates that more than half of the generated responses are discarded, highlighting the significant inefficiency in response utilization during the exploration of the diversity. In contrast, our proposed DCPO method achieves an average RUR of approximately 70% after the first epoch. This represents a substantial improvement, with an absolute increase of 28% and a relative improvement of 64% over the GRPO’s.
@@
model GRPO GSPO DCPO
Qwen2.5-Math-1.5B-Instruct 45.6% - 67.1%
Qwen2.5-3B 48.3% - 74.3%
Qwen2.5-Math-7B 37.4% 43.5% 73.2%
Qwen2.5-14B 43.9% 47.6% 72.4%
Average 43.8% 45.6% 71.8%
@@ 5.4 Ablation Studies
We conducted an ablation study on Qwen2.5-Math-7B to assess the contribution of each component in DCPO, using Avg@32 as the evaluation metric. This metric highlights the robustness and stability of the learned policy distribution. To ensure fairness, each experiment modifies a single component of the baseline GRPO framework while keeping all other settings identical and removing the KL divergence term to align with DAPO, GSPO, and the full DCPO. The learning curves among 20 to 400 steps are shown in Figure 3.
@@ A.3 Advantage Calculation
In previous works, including GRPO and DAPO, the advantage $\hat{A}_{j,t}^{i}$ for the token $t$ in the response $j$ is calculated by standardizing the reward $R^{i}_{j}$ against the mean $\mu^{i}$ and standard deviation $\sigma^{i}$ of the rewards of the $G$ responses generated in the $i$-th step, as shown in  Equation 11.
@@ A.3.1 GRPO
Shao et al. (2024) proposed the Group Relative Policy Optimization(GRPO), which first samples $G$ responses for each query, assigns rewards $R$ through a rule-based reward function and estimates token-level advantages as in Equation 11. Finally, the model parameters are updated using Equation 12, referred to as Sequence Level Mean loss (SLM) as: